Use the following information and calculate the quick ratio for Davis Company and for Bender Inc. Davis Co. Bender Inc. Account Dr. Cr. Dr. Cr. Cash $321 $425 Cash equivalents 88 95 Current notes receivable 56 46 Accounts receivable 603 307 Prepaid expenses 55 85 Merchandise inventory 714 898 Fixed assets 920 755 Accumulated depreciation—Fixed assets $415 $225 Accounts payable 260 198 Current accrued liabilities 213 149 Mortgage payable 917 824 Capital 952 1,215 Total $2,757 $2,757 $2,611 $2,611 a. Calculate the quick ratio for each company. Round your answers to two decimal places. Davis Co. Quick ratio: Bender Inc. Quick ratio: b. which one is more able to meet current liabilities. would be more liquid.
Answer #1

Answer a:

Quick ratio = (Cash + Cash equivalents + Current notes receivable + Accounts receivable) / (Accounts payable + Current accrued liabilities)

Davis Co: Quick ratio = (321 + 88 + 56 + 603) / (260 + 213) = 2.26

Bender Inc.: Quick ratio = (425 + 95 + 46 + 307) / (198 + 149) = 2.52

Hence:

Davis Co Quick ratio = 2.26

Bender Inc. Quick ratio = 2.52

Answer b:

Bender Inc.

Both Davis co and Bender Inc. have acceptable quick ratios. As Bender Inc.'s quick ratio is higher than that of Davis Co., Bender Inc. is more liquid.